### v1

Abstract
In clinical practice, medical image segmentation provides useful information on the contours and dimensions of target organs or tissues, facilitating improved diagnosis, analysis, and treatment. In the past few years, convolutional neural networks (CNNs) and Transformers have dominated this area, but they still suffer from either limited receptive fields or costly long-range modeling. Mamba, a State Space Sequence Model (SSM), recently emerged as a promising paradigm for long-range dependency modeling with linear complexity. In this paper, we introduce a Large Window-based Mamba U-shape Network, or LMa-UNet, for 2D and 3D medical image segmentation. A distinguishing feature of our LMa-UNet is its utilization of large windows, excelling in locally spatial modeling compared to small kernel-based CNNs and small window-based Transformers, while maintaining superior efficiency in global modeling compared to self-attention with quadratic complexity. Additionally, we design a novel hierarchical and bidirectional Mamba block to further enhance the global and neighborhood spatial modeling capability of Mamba. Comprehensive experiments demonstrate the effectiveness and efficiency of our method and the feasibility of using large window size to achieve large receptive fields. Codes are available at https://github.com/wjh892521292/LMa-UNet.
Keywords:
Medical Image Segmentation UNet Mamba.
1 Introduction
Efficiently segmenting biomedical objects of interest (e.g., lesions) in large-size 2D/3D images significantly enhances downstream clinical practice and biomedical research. Currently, automatic segmentation models, leveraging popular deep learning backbones such as convolutional neural networks (CNNs) and Transformers [19], have reduced manual annotation requirements but may incur considerable computational costs or overlook specific details. CNN-based models (e.g., UNet [18]) typically seek to extract global patterns by hierarchically stacking small kernels, excelling in pixel-level feature extraction but being ineffective in learning long-range dependencies due to their limited receptive fields [14]. Though recent investigations [3] have shown the effectiveness of large convolution kernels, it often requires specific optimization strategies and complicated inference-time model reformulation. In contrast, Transformer-based algorithms offer powerful long-range modeling but sacrifice pixel-level spatial modeling [21, 20]. Further, a key component, the self-attention module, incurs quadratic complexity and cannot handle excessive tokens [9], resulting in the need to pack pixels into windows and hence sacrificing resolution information. Especially, many studies have shown that Transformers achieve the best trade-off at a $7\times 7$ window size as a smaller window causes more computation and a larger window causes a significant drop in performance [13, 1]. Later studies have investigated CNN-Transformer hybrid models or approaches for intra-patch locality modeling [20, 23]. However, due to the typically larger sizes of medical images compared to natural images, how to reduce the interaction complexity between global patches and how to enlarge the receptive field of local spatial modeling are still open.
Recently, structured state space sequence models (SSMs) [6] (e.g., Mamba [5]) have emerged as a powerful long-sequence modeling approach with linear complexity in terms of input size, shedding light on efficient modeling of both local and global dependencies. In contrast to the conventional self-attention approach, SSMs enable each element in a 1D array (e.g., a text sequence) to interact with any of the previously scanned samples through a compressed hidden state, effectively reducing the quadratic complexity to linear. SSMs were devised to address natural language processing (NLP) tasks [4], but also show effectiveness in computer vision [25]. For medical image segmentation, for example, U-Mamba [5] and SegMamba [22] introduced SSM-CNN hybrid models that directly apply Mamba to efficiently model long-range dependencies at the pixel-level. Although effective, the potential and deficiency of Mamba are still not yet fully explored and resolved, in three aspects. First, benefiting from its linear complexity, Mamba possesses more flexibility in space allocation. That is, unlike the stereotypes of previous methods based on small convolution kernels or size-constrained self-attention windows, Mamba is promising in endowing the model with an ability in large receptive field spatial modeling, which has been neglected in the known Mamba-based approaches. Second, Mamba is a unidirectional sequence modeling method that lacks positional awareness and focuses more on posterior tokens. Third, the original Mamba was proposed for 1D sequence modeling, which is not suitable for computer vision tasks that require spatial-aware understanding. Mainly due to local adjacent spatial pixels becoming discontinuous, Mamba can suffer forgetting problems and inefficient local modeling capabilities.
In this paper, we propose a Large-window-based Mamba U-shape Network (LMa-UNet) for 2D and 3D medical image segmentation. LMa-UNet utilizes the powerful sequence modeling capabilities and linear complexity of Mamba to implement large receptive fields by assigning large windows to SSM modules. Further, we design a novel hierarchical and bidirectional large window-based Mamba block (LM block) to enhance the representation modeling capability of SSMs. The bidirectional Mamba design is used for location-aware sequence modeling, reducing the weight impact of the input order. The hierarchical Mamba design is composed of two types of SSM operations: (i) pixel-level SSM (PiM) and (ii) patch-level SSM (PaM). PiM captures the neighborhood and local-scope pixel information by large window SSM and avoids the forgetting problem that occurs in SSMs due to discontinuity of adjacent patches after tokenization. PaM deals with long-range dependency modeling and global patch interaction.
Our main contributions are as follows. (1) We propose a Large window-based Mamba UNet (LMa-UNet) for 2D/3D medical image segmentation. (2) We assign windows of large receptive fields to SSM layers to enable the model to possess the capability of large spatial modeling. (3) We design a bidirectional Mamba for location-aware sequence modeling. (4) We propose a novel hierarchical Mamba module composed of pixel-level SSM (PiM) and patch-level SSM (PaM), enhancing local-neighborhood pixel-level feature modeling and long-range global patch-level modeling.
2 Preliminaries: SSM Models
SSM-based models, namely the structured state space sequence models (S4) and Mamba [5], all evolved from the continuous system that maps a 1-dimensional function or sequence $x(t)\rightarrow y(t)\in\mathcal{R}$ through a hidden state $h(t)\in\mathcal{R}^{N}$. This process can be represented as the following linear ordinary differential equation:
$\begin{split}h^{\prime}(t)&=\mathbf{A}h(t)+\mathbf{B}x(t),\\ y(t)&=\mathbf{C}h(t),\end{split}$ (1)
where $\mathbf{A}\in\mathcal{R}^{N\times N}$ is a state matrix and $\mathbf{B},\mathbf{C}\in\mathcal{R}^{N}$ are projection parameters.
S4 and Mamba are discrete versions of the aforementioned continuous system, which include a timescale parameter $\mathbf{\Delta}$ to transform the continuous parameters $\mathbf{A},\mathbf{B}$ to discrete parameters $\overline{\mathbf{A}},\overline{\mathbf{B}}$. Typically, the zero-order hold (ZOH) is employed as the discretization rule and can be defined as follows:
$\begin{split}\overline{\mathbf{A}}&=\text{exp}(\Delta\mathbf{A}),\\ \overline{\mathbf{B}}&=(\mathbf{\Delta}\mathbf{A})^{-1}(\text{exp}(\mathbf{\Delta}\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}.\end{split}$ (2)
After the discretization of $\mathbf{A},\mathbf{B}$, the discretized version of Eq. (1) (i.e., SSM models) can be written as:
$\begin{split}h^{\prime}(t)&=\overline{\mathbf{A}}h(t)+\overline{\mathbf{B}}x(t),\\ y(t)&=\mathbf{C}h(t).\end{split}$ (3)
Then the output is computed through a global convolution, defined as:
$\begin{split}\overline{\mathbf{K}}&=(\mathbf{C}\overline{\mathbf{B}},\mathbf{C}\overline{\mathbf{AB}},\mathbf{C}\overline{\mathbf{A}}^{L-1}\overline{\mathbf{B}}),\\ y&=x*\overline{\mathbf{K}},\end{split}$ (4)
where $L$ is the length of the input sequence $x$ and $\overline{\mathbf{K}}\in\mathcal{R}^{L}$ is a structured convolutional kernel.
3 Method
In this section, we first introduce the overall architecture of LMa-UNet. Subsequently, we elaborate the core component, the LM block.
Figure 1: An overview of our proposed LMa-UNet.
3.1 LMa-UNet
An overview of LMa-UNet is given in Fig. 1. Specifically, except for the common UNet composition of a depth-wise convolution, an encoder with downsampling layers, a decoder with upsampling layers, and skip-connection, LMa-UNet improves the structure of UNet with proposed large window-based Mamba (LM) blocks inserted into the encoder. Given a 3D input image with a resolution of $C\times D\times H\times W$, the depth-wise convolution first encodes the input into a feature map $\mathbf{F}_{0}\in\mathcal{R}^{48\times\frac{D}{2}\times\frac{H}{2}\times\frac{W}{2}}$. Then the feature map $\mathbf{F}_{0}$ is fed into each LM block and the corresponding down-sampling layers, and multi-scale feature maps are obtained. An LM block contains two Mamba modules: pixel-level SSM (PiM) and patch-level SSM (PaM). For the $l^{th}$ layer, the process can be formulated as:
$\mathbf{F^{\prime}}_{l}=\text{PiM}(\mathbf{F}_{l}),\quad\mathbf{F^{\prime\prime}}_{l}=\text{PaM}(\mathbf{F^{\prime}}_{l}),\quad\mathbf{F}_{l+1}=\text{Down-sampling}(\mathbf{F^{\prime\prime}}_{l}),$ (5)
where PiM and PaM denote pixel-level SSM and patch-level SSM, respectively. Down-sampling denotes the down-sampling layers. After each stage, the produced feature map $\mathbf{F}_{l+1}$ is encoded to $(2C_{l},\frac{D_{l}}{2},\frac{H_{l}}{2},\frac{W_{l}}{2}$), where $C_{l},(D_{l},H_{l},W_{l})$ represent the channel and resolution of feature map $\mathbf{F}_{l}$. As for the decoder part, we adopt a UNet decoder and residual block with skip connections for upsampling and predicting the final segmentation masks.
3.2 LM block
LM block is our core component used for further spatial modeling of feature maps of different scales at each stage. Different from the previous methods that use CNN for local pixel-level modeling and Transformer for long-range patch-level dependency modeling, an LM block can accomplish pixel-level and patch-level modeling simultaneously, benefiting from the linear complexity of Mamba. More critically, a lower complexity allows setting larger windows to obtain bigger receptive fields, which will improve the efficiency of local modeling, as shown in Fig. 2(a). Specifically, LM block is a hierarchical design that consists of pixel-level SSM (PiM) and patch-level SSM (PaM); the former is used for local-neighborhood pixel modeling and the latter is used for global long-range dependency modeling. Besides, each Mamba layer in an LM block is bidirectional, which is proposed for location-aware sequence modeling.
Pixel-level SSM (PiM). Since Mamba is a continuous model, the discrete nature of input pixels can weaken the correlation modeling of locally adjacent pixels. Hence, we propose a pixel-level SSM to split the feature map into multiple large sub-windows and perform SSM operations on the sub-windows. We first equally divide a whole feature map into non-overlapping sub-windows for 2D or sub-cubes for 3D. Take 2D for example. Given an input of $H\times W$ resolution, we divide the feature map into sub-windows of size $m\times n$ each ($m$ and $n$ can be up to 40). Without loss of generality, we assume that $H/m$ and $W/n$ are both integers. Then we have $\frac{HW}{mn}$ sub-windows, as shown in Pixel-level SSM of Fig. 1. Under this scheme, when these sub-windows are sent into a Mamba layer, the local adjacent pixels will be input continuously into SSM; thus, the relation between local-neighborhood pixels can be better modeled. Moreover, under the large window partition strategy, the receptive fields are enlarged and the model can obtain more details of the local pixels. However, the image is divided into non-overlapping sub-windows. Hence, we need a mechanism for communication between different sub-windows, for long-range dependency modeling.
Figure 2: (a) Respective field comparison among CNN, Transformer, and our proposed LMa-UNet. CNNs often use small kernels (like $3\times 3$), and Transformers often use $7\times 7$ sized windows. (b) Scanning order comparison of vanilla Mamba vs. our proposed bidirectional Mamba.
Patch-level SSM (PaM). We introduce a patch-level SSM layer to impart information among different sub-windows. As shown in Patch-level SSM of Fig. 1, a feature map $\mathbf{F^{\prime}}_{l}$ of resolution $H\times W$ first passes through a pooling layer of size $m\times n$ to allow important information for each of the $\frac{HW}{mn}$ sub-windows to be summarized into a single representative. Thus, we obtain aggregate maps $\mathbf{Z}_{l}$ with $\frac{HW}{mn}$ representatives, which are then used to communicate among the sub-windows through Mamba for global-range dependency modeling. After the interaction in Mamba, we unpooling the aggregate maps back to the same size as the initial feature map $\mathbf{F^{\prime}}_{l}$, and apply a residual connection. The process of $\mathbf{F}^{\prime\prime}_{l}=\text{PaM}(\mathbf{F}^{\prime}_{l})$ in Eq. (5) can be carried out as:
$\mathbf{W}_{l}=\text{Pooling}(F^{\prime}_{l}),\quad\mathbf{W}^{\prime}_{l}=\text{Bi-Mamba}(\mathbf{W}_{l}),\quad\mathbf{F^{\prime\prime}}_{l}=\text{Unpooling}(\mathbf{W}^{\prime}_{l}),$ (6)
where Pooling and Unpooling denote the pooling layer and unpooling layer, respectively. Bi-Mamba denotes the proposed bidirectional Mamba layer.
Bidirectional Mamba (BiM). Different from the vanilla Mamba block which is based on forward-only scan direction SSM layers, each SSM layer (including PiM and PaM) in our LM block is bidirectional. Fig. 2(b) shows the differences. In the original Mamba, as a continuous model, some information forgetting occurs on the elements entered earlier, and the latest elements that enter into Mamba will retain much more information. Thus, the original Mamba with a single scanning direction will focus more on the posterior patches, rather than the center areas of the feature maps often with more organs and lesions. To this end, we propose a bidirectional Mamba structure by performing both forward and backward scanning at the same time and superimposing the output results. The detailed structure is shown in the left part of Fig. 1. BiM has two advantages. First, the model can focus more on the informative patches in the center areas of the image likely with more organs and lesions, rather than the corner areas. Second, for each patch, the absolute position information and relative position information with other patches can be well modeled by the network.
4 Experiments
4.1 Datasets
We conduct experimental comparisons with state-of-the-art methods on two datasets for 2D and 3D segmentation tasks to validate the effectiveness and scalability of LMa-UNet.
Abdomen CT. Abdomen CT is a publicly available 3D multi-organ segmentation dataset comprising 100 CT cases from the MICCAI 2022 FLARE Challenge [16], including 13 types of abdominal organs (liver, spleen, pancreas, right kidney, left kidney, stomach, gallbladder, esophagus, aorta, inferior vena cava, right adrenal gland, left adrenal gland, and duodenum). The size of a 3D CT image is 40 $\times$ 224 $\times$ 192. 50 cases from the MSD Pancreas dataset with annotations from AbdomenCT-1K are used for training, and another 50 cases from different medical centers [2] are used for testing.
Abdomen MR. Abdomen MR is a publicly available 2D segmentation dataset comprising 110 MRI cases from the MICCAI 2022 AMOS Challenge [11], including 13 types of abdominal organs (the same as the Abdomen CT dataset). The size of a 2D MRI image is 320 $\times$ 320. Following the previous work [15, 22], 60 annotated cases are used for training, and another 50 cases are used for testing.
Table 1: Quantitative segmentation results on the 3D Abdomen CT dataset and 2D Abdomem MR dataset.
  3D Abdomen CT   2D Abdomen MR
Method  DSC $\uparrow$  NSD $\uparrow$  DSC $\uparrow$  NSD $\uparrow$
SegResNet [17] 79.27 82.57 73.17 80.34
nnUNet [10] 86.15 89.72 74.50 81.53
UNTER [8] 68.24 70.04 57.47 63.09
SwinUNTER [7] 75.94 76.63 70.28 76.69
nnFormer [24] 78.34 81.45 72.79 79.63
U-Mamba (concurrent) [15] 86.38 89.80 76.25 83.27
LMa-UNet(Ours) 86.82 90.02 77.35 83.80
4.2 Implementation Setup
Our LMa-UNet is implemented on PyTorch 1.9.0 based on the nnU-Net framework. All the experiments are conducted on an NVIDIA GeForce RTX 3090 GPU. The batch size in training is 2 for the 3D dataset (Abdomen CT) and 24 for the 2D dataset (Abdomen MR). The Adam [12] optimizer with momentum = 0.99 is used. The initial learning rate is 0.01 with a weight decay of 3e-5. The maximum training epoch number is 1000. For the Abdomen CT dataset, the stage is 6 but the dimensions are not consistent; thus we set the rectangle window size with three dimensions to [20, 28, 24], [20, 28, 24], [10, 14, 12], [10, 14, 12], [5, 7, 6], and [5, 7, 6] for each stage. For the Abdomen MR dataset, the stage is 7 and the window size is 40, 20, 20, 10, 10, 5, and 5 for each stage.
4.3 Overall Performances
The baseline models include three types of representative medical image segmentation networks: CNN-based networks (nnU-Net [10] and SegResNet [17]), Transformer-based networks (UNETR [8], SwinUNETR [7], and nnFormer [24]), and the latest Mamba-based network (U-Mamba [15]). For fair comparison, we also implement all the models in the nnU-Net framework, and use the default image pre-processing. Table 1 presents the results. Compared to both CNN-based and Transformer-based segmentation methods, our proposed LMa-UNet achieves improved performances in both DSC and NSD, which indicates that the global modeling capabilities of Mamba are critical to medical image segmentation. Note that, compared to U-Mamba which simply applies Mamba as a global modeling adapter, LMa-UNet exhibits improvements over U-Mamba, validating the effectiveness of our bidirectional and hierarchical Mamba designs. These results also demonstrate the potential of Mamba in global and local feature modeling with large window-based receptive fields.
Table 2: Performances of LMa-UNet in three different window size settings. The window size sequence indicates the window size in each stage.
Window size [10, 5, 5, 5, 5, 5, 5] [20, 10, 10, 10, 5, 5, 5] [40, 20, 20, 10, 10, 5, 5]
DSC 75.89 76.45 77.35
NSD 82.26 82.78 83.80
Table 3: Performances of LMa-UNet with different sub-modules. PiM = Pixel-level SSM. PaM = Patch-level SSM. BiM = Bidirectional Mamba.
Method Baseline Only PiM Only PaM PiM + BiM PaM + BiM PiM + PaM PiM + PaM + BiM
DSC 74.50 76.82 76.22 76.90 76.73 77.10 77.35
NSD 81.53 83.05 82.59 83.31 82.94 83.54 83.80
4.4 Is the Window Size of LMa-UNet Important?
Due to the linear complexity of Mamba, LMa-UNet can achieve large window modeling that CNNs and Transformers cannot. To validate Mamba’s large spatial modeling potential, we explore LMa-UNet’s performance in different window size settings. Table 2 shows the results on the Abdomen MR dataset, whose stage number is 7. Comparing the performances of the three window-size settings, one can see that LMa-UNet with larger window sizes achieves better performances. This indicates that large receptive fields are critical for medical image segmentation which can be achieved with Mamba due to its linear complexity. This will provide more possibilities for future large kernel or window modeling.
4.5 Ablation Study
To validate the effect of each key component in our LMa-UNet, we conduct ablation experiments on the Abdomen MR dataset. Table 3 shows the results. One can see that both PiM and PaM provide improvements for LMa-UNet over the baseline model, validating the effectiveness of PiM and PaM in local pixel-level modeling and global modeling, respectively. Notably, the model with PiM gains more improvements than that with PaM, suggesting that enlarging the receptive field of local feature modeling is a key to improving model performance. After introducing BiM, the performance of LMa-UNet further improves, which shows the importance of bidirectional Mamba for location-aware sequence modeling. Finally, LMa-UNet with all the components achieves the best performance, which further demonstrates the effectiveness of our method and its components.
5 Conclusions
In this paper, we introduced a new Mamba-based UNet model for medical image segmentation, achieving large window spatial modeling. Further, we designed a bidirectional and hierarchical SSM to enhance the capacities of Mamba in local and global feature modeling. Comprehensive experiments on multi-organ segmentation datasets demonstrated the effectiveness of our proposed method. In the future, we will explore the extensibility and applications of SSMs in other medical imaging tasks, such as detection, registration, reconstruction, etc.